Scores of ten students in a spelling test.

11, 8, 12, 7, 5, 3, 4, 15, 9, 6. Find the variance and standard deviation of ungrouped data.


Answers:



1. Variance =



2. Standard Deviation =


1
Expert's answer
2022-01-10T12:28:52-0500

1.

Variance:

"\sigma^2=\frac{\sum (x_i-\overline{x})^2}{n}=13.00"

where

mean:

"\overline{x}=\frac{\sum x_i}{n}=8"

"n=10"


2.

standard deviation:

"\sigma=\sqrt{13}=3.60"